OpenStudy (anonymous):

Use implicit differentiation to find ∂z/∂x and ∂z/∂y. x^3 + y^3  +z^3 + 6xyz=  1

OpenStudy (anonymous):

does any one remember

ganeshie8 (ganeshie8):

when you differentiate with.respect.to "x", y an z are treated as contants

OpenStudy (anonymous):

yes but its respect as function of z

ganeshie8 (ganeshie8):

\(\large x^3 + y^3 +z^3 + 6xyz= 1\) \(\large \frac{\partial }{\partial x} (x^3 + y^3 +z^3 + 6xyz= 1)\)

OpenStudy (anonymous):

so do i take both der of x and also to z so do i chain the second one at z

ganeshie8 (ganeshie8):

lets see

OpenStudy (anonymous):

z is defined implicitly as a function of

OpenStudy (anonymous):

x

ganeshie8 (ganeshie8):

\(\large \frac{\partial }{\partial x} (x^3 + y^3 +z^3 + 6xyz= 1)\) \(\large 3x^2 + 3z^2 \frac{\partial z}{\partial x} + 6y( x\frac{\partial z}{\partial x} + z)= 0\)

OpenStudy (anonymous):

i kn 3x^2+3z^2dz/dx+6yz(dz/dx)+6xy(dz/dx)=0

ganeshie8 (ganeshie8):

you can solve \(\frac{\partial z}{\partial x}\) from above

ganeshie8 (ganeshie8):

you may also first solve \(z\) explicitly, and then find the partial and verify if you get the same

ganeshie8 (ganeshie8):

nvm, we cannot solve \(z\) explicitly

OpenStudy (anonymous):

i was about to ask you to show me =P

OpenStudy (anonymous):

okay thx

ganeshie8 (ganeshie8):

lol

ganeshie8 (ganeshie8):

u wlc :)